# DisCo: Disentangled Control for Referring Human Dance Generation in Real World

### v1

@@ Abstract
Generative AI has made significant strides in computer vision, particularly in image/video synthesis conditioned on text descriptions. Despite the advancements, it remains challenging especially in the generation of human-centric content such as dance synthesis. Existing dance synthesis methods struggle with the gap between synthesized content and real-world dance scenarios. In this paper, we define a new problem setting: Referring Human Dance Generation, which focuses on real-world dance scenarios with three important properties: (i) Faithfulness: the synthesis should retain the appearance of both human subject foreground and background from the reference image, and precisely follow the target pose; (ii) Generalizability: the model should generalize to unseen human subjects, backgrounds, and poses; (iii) Compositionality: it should allow for composition of seen/unseen subjects, backgrounds, and poses from different sources. To address these challenges, we introduce a novel approach, DisCo, which includes a novel model architecture with disentangled control to improve the faithfulness and compositionality of dance synthesis, and an effective human attribute pre-training for better generalizability to unseen humans. Extensive qualitative and quantitative results demonstrate that DisCo can generate high-quality human dance images and videos with diverse appearances and flexible motions. $*{}^{*}$qual Contribution    ${}^{\mathsection}$ork done during internship at Microsoft
@@ 1 Introduction
Generative AI has garnered significant interests in computer vision community. Recent advancements in text-driven image and video synthesis (T2I/T2V) [47, 44, 20, 51, 72, 18, 57, 9, 71] bolstered by the advent of diffusion models [19, 8, 59, 60] exhibited remarkable ingenuity and generative quality, demonstrating considerable potential in image and video synthesis, editing, and animation. However, the synthesized images/videos are still not satisfying, especially for human-centric applications, e.g., human dance synthesis.
Despite the long history of human dance synthesis, existing methods greatly suffer from the gap between the synthesized content and real-world dance scenarios. Starting from the era of GAN [14, 3, 27], researchers [67, 66, 11, 6] try to extend the video2video style transfer for transferring dance movements from a source video to a target individual, which often requires human-specific fine-tuning on the target person. Recently, a line of work [29, 39] leverages pre-trained diffusion-based T2I/T2V models to generate dance images/videos conditioned on text prompts. Such coarse-grained condition greatly limits the degree of the controllability, making it almost impossible for users to precisely specify the anticipated subjects (i.e., human appearance) as well as the dance-moves (i.e., human pose). Though the introduction of ControlNet [73] partially alleviates this problem by incorporating pose control with geometric human keypoints, it remains unclear how ControlNet can ensure the consistency of rich semantics (such as human appearance) in the reference image, due to its dependency on text prompt. Moreover, almost all existing methods trained on limited dance video datasets suffer from either limited subject attributes [67, 6, 77] or excessively simplistic scenes and backgrounds [34, 38, 26, 25], leading to the poor zero-shot generalizability to unseen composition of human subjects, poses and backgrounds.
(a) Faithfulness
(b) Generalizability
(c) Compositionality
Figure 1: We propose DisCo for referring human dance generation, which can generate human dance images/videos with the following three properties: (a) Faithfulness: retaining the appearance of foreground (FG) and background (BG) in consistent to the reference image while precisely following the pose; (b) Generalizability: generalizable to unseen human subject FG, BG and pose; (c) Compositionality: adapting to arbitrary composition of human subject FG, background and pose, each from a different source.
In order to support real-life applications, such as user-specific short video content generation, we define a new problem setting: referring human dance generation, with the focus on real-world dance scenarios. Given a reference image designating a human subject foreground and the background together (or two images depicting the human foreground and background respectively), and the dance movement specified by a pose or a sequence of poses, a model should be able to synthesize human dance images/videos with the following three properties:
  • Faithfulness: The generated images/videos should retain the appearance of human subjects and backgrounds, in consistent to the reference images (Figure 1a). In addition, the generated human subject should precisely follow the provided pose.
@@
  • Generalizability: The model should be able to generalize to unseen human subject, background and pose (Figure 1b), without the need of human-specific fine-tuning.
@@
  • Compositionality: The generated images/videos can be from an arbitrary composition of seen or unseen human subject, background and pose, sourced from different images/videos (Figure 1c).
@@
In this regard, we propose a novel approach, DisCo, for referring human dance generation in real world. DisCo consists of two key designs: ($i$) a novel model architecture with disentangled control for improved faithfulness and compositionality; and ($ii$) an effective pre-training strategy with DisCo for better generalizability, named human attribute pre-training.
Model Architecture with Disentangled Control (Section 3.2): We propose an organic integration of conditions with cross-attention and ControlNet. Specifically, we substitute the text condition in T2I diffusion model with the CLIP image embeddings of the human subject, which is incorporated via the cross-attention modules of U-Net; while the background and human pose conditions are fed into two separate ControlNet branches. By disentangling the control from all three conditions, DisCo can not only achieve fidelity in human foregrounds and backgrounds but also enable arbitrary compositionality of human subjects, backgrounds, and dance-moves (as shown in Figure 1).
@@
Human Attribute Pre-training (Section 3.3): We design a proxy task in which the model conditions on the separate foreground and background areas and must reconstruct the complete image. In this way, the model learns to better encode-and-decode the complicated human faces and clothes during pre-training, and leaves the pose control learning to the fine-tuning stage of human dance synthesis. Crucially, without the constraint of pairwise human images for pose control, we can leverage large-scale collections of human images to learn diverse human attributes, in turn, greatly improve the generalizability of DisCo to unseen humans.
@@
  • We define a new problem setting of generating real-world dance content, referring human dance generation, to facilitate its potential application in the production of user-specific short videos.
  • To address this problem, we propose DisCo framework with ($i$) a novel model architecture for disentangled control to ensure faithfulness and compositionality in generation; and ($ii$) human attribute pre-training to improve generalizability to unseen humans.
  • We conduct a broad variety of evaluations and applications to demonstrate the effectiveness of DisCo, including human image editing and human dance video synthesis.
@@ 2 Related Work
Different from GANs [14, 3, 27], recent diffusion probabilistic models [58, 19, 8, 59, 60] trained with totally new objectives and pipeline have shown great success in high-quality image/video generation. Towards user-specific generation, text prompts are first utilized as the condition for image generation [47, 44, 20, 51, 72]. Among these methods, Stable Diffusion [48] (SD) stands as the representative work to date, with high efficiency and competitive quality via diffusion over the latent space. For better controllability, ControlNet [73] introduces additional control mechanisms into SD beyond texts, such as sketch, human skeleton, and segmentation map. Compared to text-to-image synthesis, text-to-video synthesis [18, 57, 9, 71, 29], is relative new and still remains challenging, due to ($i$) lack of well-annotated video-text data, hence expensive to scale; ($ii$) difficulties in modeling temporal consistency, to ensure realistic actions, appearance and scenes; and ($iii$) lack of compositional understanding for human-object or object-object interaction generation. Thus, controllable video generation methods stem from pre-trained text-to-image models, especially SD. Follow-your-pose [39] tries to introduce pose condition into video generation with image-pose pairs and pose-free videos. Text2Video-Zero [29] propose to modify the pre-trained T2I model to enable the zero-shot human video generation. In this work, we look into a more challenging setting of conditional human image/video synthesis, specifically referring human dance generation, which requires precise control of both human attributes (such as identity, clothing, makeup, hairstyle, etc.) and the dance-moves (poses).
@@
Early work on this task includes video-to-video synthesis [67, 66, 11, 6], still image animation [56, 70, 21, 55, 68, 75, 2, 40] and motion transfer [77, 54, 33, 61, 12]. Nevertheless, these lines of methods require either a several-minute-long target person video for human-specific fine-tuning, or multiple separate networks and cascaded training stages for background prediction, motion representation and occlusion map generation. The advances of diffusion models [48] greatly simplify the training of such generative models, inspiring follow-up diffusion models tailored for human dance generation. For example, [31] augments the video generation process with a set of user-specified strokes, and [43] synthesizes an optical flow sequence with the given text prompt for action generation. However, these methods require a separate motion prediction module and struggle to precisely control the human pose. DreamPose [26] is perhaps the most relevant study to ours, which proposes an image-and-pose conditioned diffusion method for still fashion image animation. However, as they consider only fashion subjects with easy catwalk poses in front of an empty background, their model may suffer from limited generalization ability, prohibiting its potential for more intricate human dance synthesis in real-world scenarios.
@@ 3 DisCo
We start by first formally define the problem setting for referring human dance generation. Let $f$ and $g$ represent human foreground and background in the reference image. Given a specific (or a sequence of) pose keypoint $p=p_{t}$ (or $p=\{p_{1},p_{2},...,p_{T}\}$), we aim to generate realistic images $I_{t}$ (or videos $V=\{I_{1},I_{2},...,I_{T}\}$) conditioned on $f,g,p$. The generated images (or videos) should be 1) faithful: the human attribute and background of the synthesis should be consistent with $f$ and $g$ from the reference image and the generated human subject should be aligned with the pose $p$; 2) generalizable: the model should be able to generalize to unseen humans, backgrounds and poses, without the need of human-specific fine-tuning; and 3) composable: the model should adapt to arbitrary composition of $f,g,p$ from different image/video sources to generate novel images/videos. In what follows, Section 3.1 briefly reviews the latent diffusion models and ControlNet, which are the basis of DisCo. Section 3.2 details the model architecture of DisCo with disentangled control of human foreground, background, and pose to enable faithful and fully composable human dance image/video synthesis. Section 3.3 presents how to further enhance the generalizability of DisCo, as well as the faithfulness in generated contents by pre-training human attributes from large-scale human images. The overview of DisCo can be found in Figure 2.
3.1 Preliminary: Latent Diffusion Models & ControlNet
@@
Latent Diffusion Models (LDM) is a type of diffusion model that operates in the encoded latent space of an autoencoder $\mathcal{D}(\mathcal{E}(\cdot))$, therefore saving the computational and time complexity. An exemplary LDM is the popular Stable Diffusion (SD) [48] which consists an autoencoder VQ-VAE [65] and a time-conditioned U-Net [49] for noise estimation. Since SD is designed for T2I, a CLIP ViT-L/14 text encoder [46] is used to project the input text query into the text embedding condition $c_{\text{text}}$.
During training, given an image $I$ and the text condition $c_{\text{text}}$, the image latent $z_{0}=\mathcal{E}(I)$ is diffused in $T$ time steps with a deterministic Gaussian process to produce the noisy latent $z_{T}\sim\mathcal{N}(0,1)$. SD is trained to learn the reverse denoising process with the following objective [48]:
@@
$L=\mathbb{E}_{\mathcal{E}(I),c_{\text{text}},\epsilon\sim\mathcal{N}(0,1),t}\left[\lVert\epsilon-\epsilon_{\theta}(z_{t},t,c_{\text{text}})\rVert_{2}^{2}\right],t=1,...,T$
@@
where $\epsilon_{\theta}$ represents the trainable modules, containing a U-Net architecture composed of the convolution (ResBlock) and self-/cross-attention (TransBlock), which accepts the noisy latents $z_{t}$ and the text embedding condition $c_{\text{text}}$ as the input. After training, one can apply a deterministic sampling process (e.g., DDIM [59]) to generate $z_{0}$ and pass it to the decoder $\mathcal{D}$ towards the final image.
@@
ControlNet [73], built upon SD, manipulates the input to the intermediate layers of the U-Net in SD so as to further control the overall behavior of SD. Specifically, it creates a trainable copy of the U-Net down/middle blocks and adds an additional “zero convolution” layer, i.e., $1\times 1$ convolution layer with both weight and bias initialized with zeros. The outputs of each copy block is then added to the skip connections of the original U-Net. Apart from the text condition $c_{\text{text}}$ captured by cross-attention modules in SD, ControlNet is trained with an additional external condition vector $c$ which can be many types of condition, such as edge map, pose, depth map and segmentation map. Training on a specific domain of $c$ results in an effective conditional T2I generation.
@@ 3.2 Model Architecture with Disentangled Control
The direct application of ControlNet to referring human dance generation presents challenges due to the missing of reference human image condition, which is critical for keeping the human identity and attribute consistent in the synthesized images. Recent explorations in image variations [32] replace the CLIP text embedding with the CLIP image embedding as the SD condition, which can retain some high-level semantics from the reference image. Nevertheless, the geometric/structural control onto the generated image is still missing.
(a) Model Architecture with Disentangled Control
(b) Human Attribute Pre-training
@@
Figure 2: The proposed DisCo framework for referring human dance generation.
@@
Taking the distinctive benefits of these two different control designs, we introduce a novel model architecture with disentangled control, to enable accurate alterations to the human pose, while simultaneously maintaining attribute and background stability. Meanwhile, it also facilitates full compositionality in the human dance synthesis, accommodating any combination of human foreground, pose, and background (Figure 1c). Specifically, given a reference human image, we can first utilize an existing human matting method (e.g., SAM [30, 37]) to separate the human foreground from the background. Next, we explain how all three conditions, the human foreground $f$, the background $g$ and the desired pose $p$, are incorporated into DisCo.
@@
Referring Foreground via Cross Attention. To help model easily adapt to the CLIP image feature space, we first use the pre-trained image variation latent diffusion model [32] for the U-Net parameter initialization. However, in contrast to using the global CLIP image embeddings employed by image variation methods, here we adopt the local CLIP image embeddings right before the global pooling layer, for more fine-grained human semantics encoding. Consequently, the original text embedding $c_{\text{text}}\in\mathbb{R}^{l\times d}$ is superseded by the local CLIP image embeddings of the human foreground $c_{f}\in\mathbb{R}^{hw\times d}$ to serve as the key and value feature in cross-attention layer, where $l,h,w,d$ represent the caption length, the height, width of the visual feature map and the feature dimension.
@@
Controlling Background and Pose via ControlNets. For pose $p$, we adopt the vanilla design of ControlNet. Specifically, we embed the pose image into the same latent space as the Unet input via with four convolution layers, and dedicate a ControlNet branch $\tau_{\theta}$ to learn the pose control. For background $g$, we insert another ControlNet branch $\mu_{\theta}$ to the model. Notably, we propose to use the pre-trained VQ-VAE encoder $\mathcal{E}$ in SD, instead of four randomly initialized convolution layers, to convert the background image into dense feature maps to preserve intricate details. The remainder of the architecture for the background ContorlNet branch follows the original ControlNet. As we replace the text condition with referring foreground in the cross-attention modules, we also update the condition input to ControlNet as the local CLIP image feature of the referring foreground. As shown in Figure 2a, the outputs of the two ControlNet branches are combined via addition and fed into the middle and up block of the U-Net.
@@
With the design of the disentangled controls above, we fine-tune DisCo with the same latent diffusion modeling objective [48]:
@@
$L=\mathbb{E}_{\mathcal{E}(I),c_{f},\tau_{\theta}(p),\mu_{\theta}(g),\epsilon\sim\mathcal{N}(0,1),t}\left[\lVert\epsilon-\epsilon_{\theta}(z_{t},t,c_{f},\tau_{\theta}(p),\mu_{\theta}(g))\rVert_{2}^{2}\right],$
@@
where $\epsilon_{\theta}$, $\tau_{\theta}$ and ${\mu_{\theta}}$ are the trainable network modules. Specifically, $\epsilon_{\theta}$ contains the U-Net architecture composed of the convolution (ResBlock) and self-/cross-attention (TransBlock), which accepts the noisy latents $z_{t}$ and the referring foreground condition $c_{f}$ as the inputs. $\tau_{\theta}$ and ${\mu_{\theta}}$ represent the two ControlNet branches for pose condition $p$ and background condition $g$, respectively.

### v3

# DisCo: Disentangled Control for Realistic Human Dance Generation

@@ Abstract
Generative AI has made significant strides in computer vision, particularly in text-driven image/video synthesis (T2I/T2V). Despite the notable advancements, it remains challenging in human-centric content synthesis such as realistic dance generation. Current methodologies, primarily tailored for human motion transfer, encounter difficulties when confronted with real-world dance scenarios (e.g., social media dance), which require to generalize across a wide spectrum of poses and intricate human details. In this paper, we depart from the traditional paradigm of human motion transfer and emphasize two additional critical attributes for the synthesis of human dance content in social media contexts: (i) Generalizability: the model should be able to generalize beyond generic human viewpoints as well as unseen human subjects, backgrounds, and poses; (ii) Compositionality: it should allow for the seamless composition of seen/unseen subjects, backgrounds, and poses from different sources. To address these challenges, we introduce DisCo, which includes a novel model architecture with disentangled control to improve the compositionality of dance synthesis, and an effective human attribute pre-training for better generalizability to unseen humans. Extensive qualitative and quantitative results demonstrate that DisCo can generate high-quality human dance images and videos with diverse appearances and flexible motions. Code is available at https://disco-dance.github.io/.
@@ 1 Introduction
Starting from the era of GAN [3, 12], researchers [52, 53] have tried to explore the human motion transfer by transferring talking and Tai-Chi poses from a source image to a target individual. This requires the generated images/videos to precisely follow the source pose and retain the appearance of human subjects and backgrounds from the target image. However, when it comes to much more diverse and nuanced visual contents such as TikTok dancing videos, GANs tend to struggle due to mode collapse, capturing only a limited portion of the real data distribution (Figure LABEL:fig:intro_figa, MRAA [53]).
Recently, diffusion-based generative models [17, 56, 57] have significantly improved the synthesis in both diversity and stability. The introduction of ControlNet [71] further enhances the controllability by injecting geometric conditions (e.g., human skeleton) into Stable Diffusion (SD) model [45], thus becomes possible to be utilized in human dance generation. However, prevailing ControlNet-based methods either rely on guidance from coarse-grained text prompts (Figure LABEL:fig:intro_figa, T2I Adaptor [40]) or simply substitute the text condition in T2I/T2V models with the referring image (Figure LABEL:fig:intro_figa, DreamPose [25]). It remains unclear how to ensure the consistency of rich human semantics and background in real-world dance scenarios. Moreover, almost all existing methods are trained on insufficient dance video data, hence suffer from either limited human attributes [63, 5, 75] or excessively simple poses and scenes [31, 36], leading to the poor zero-shot generalizability to unseen dance scenarios.
In order to support real-life applications, such as user-specific short video generation, we step from the conventional human motion transfer to realistic human dance synthesis and further highlight two properties:
@@
  • Generalizability: The model should be able to generalize to hard cases, e.g., non-generic human view as well as unseen human subject, background and pose (Figure LABEL:fig:intro_figa).
@@
  • Compositionality: The generated images/videos can be from an arbitrary composition of seen or unseen human subject, background and pose, sourced from different images/videos (Figure LABEL:fig:intro_figb).
@@
In this regard, we introduce TikTok data as the testbed and propose a novel approach, DisCo, for realistic human dance generation in social media scenarios. DisCo consists of two key designs: ($i$) a novel model architecture with disentangled control for improved compositionality; and ($ii$) an effective pre-training strategy with DisCo for better generalizability, named human attribute pre-training.
Model Architecture with Disentangled Control (Section 3.2). We attribute the failure of existing ControlNet-based methods to the inappropriate integration of various conditions. In this paper, Instead of naively combining multiple ControlNet, we utilize the VAE as the background encoder to fully leverage the prior of semantically rich images, while a tiny convolutional encoder for highly abstract skeleton. For the human subject, we incorporate its CLIP image embedding with the denoising U-Net as well as all other conditions via the cross-attention modules, to help dynamic foreground synthesis. By disentangling the control of three conditions, DisCo can not only enable arbitrary compositionality of human subjects, backgrounds, and dance-moves (Figure LABEL:fig:intro_figb), but also achieve high fidelity via the thorough utilization of the various input conditions (check Table 36 for the ablation of the condition mechanism).
@@
Human Attribute Pre-training (Section 3.3). More importantly, we design a novel proxy task in which the model conditions on the separate foreground and background region features and must reconstruct the complete image. This task is non-trivial as it enables the model to (a) effectively distinguish the dynamic human subject and static background for the ease of following pose transfer; (b) better encode-and-decode the complicated human faces and clothes during pre-training, and leaves the pose control learning to the fine-tuning stage of human dance synthesis. Crucially, without the constraint of pairwise human images for pose control, we can overcome the insufficiency of high-quality dance video data by leveraging large-scale collections of human images to learn diverse human attributes, in turn, greatly improve the generalizability of DisCo to unseen humans.
@@
  • We highlight two properties, generalizability and compositionality, that are missing from the conventional human motion transfer by introducing a more challenging social media dance synthesis problem, to facilitate its potential in the production of user-specific short videos.
  • To address these problems, we propose DisCo framework with ($i$) a novel model architecture for disentangled control to ensure compositionality in generation; and ($ii$) an effective human attribute pre-training to improve generalizability to unseen humans and non-generic views.
  • We conduct a broad variety of evaluations together with a user study, to demonstrate the effectiveness of DisCo. Notably, even without temporal consistency modeling, DisCo can already achieve superior FID (28.31 v.s 53.78) and FID-VID (55.17 v.s 66.36) scores over the state-of-the-art approaches. Adding temporal modeling further boosts FID-VID scores of DisCo to 29.37.
@@ 2 Related Work
Diffusion probabilistic models [55, 7] have shown great success in high-quality image/video generation. Towards user-specific generation, text prompts are first utilized as the condition for image generation [44, 16, 48, 69]. Among them, Stable Diffusion [45] (SD) stands as the representative work to date, with high efficiency and competitive quality via diffusion over the latent space. For better controllability, ControlNet [71] introduces additional control mechanisms into SD beyond texts, such as sketch, human skeleton, and segmentation map. Compared to image, text-to-video synthesis [18, 54, 8, 68, 27], is more challenging due to the lack of well-annotated data and difficulties in temporal modeling. Thus, existing controllable video generation methods [37, 27] mainly stem from pre-trained text-to-image model and try to introduce motion/pose prior to the text-to-video synthesis. In this work, we look into a more challenging setting of conditional human image/video synthesis which requires precise control of both human attributes (such as identity, clothing, hairstyle, etc.) as well as the dance-moves (poses) in social media dance scenarios.
@@
Early work on this task includes video-to-video synthesis [63, 64, 10, 5], still image animation [67, 19, 65, 2, 70] and motion transfer [75, 51, 30, 58]. Nevertheless, these methods require either a several-minute-long target person video for human-specific fine-tuning, or multiple separate networks and cascaded training stages for background, motion and occlusion map prediction. The advances of diffusion models [45] greatly simplify the training of such generative models, inspiring follow-up diffusion models [29, 41] tailored for human dance generation. Still, these methods require a separate motion prediction module and struggle to precisely control the human pose. DreamPose [25] is perhaps the most relevant study to ours, which proposes an image-and-pose conditioned diffusion method for still fashion image animation. However, as they consider only fashion subjects with easy catwalk poses in front of an empty background, their model may suffer from limited generalization ability, prohibiting its potential for more intricate human dance synthesis in real-world scenarios.
@@ 3 DisCo
We start by first formally introducing the social media human dance generation setting. Let $f$ and $g$ represent human foreground and background in the reference image. Given a specific (or a sequence of) pose keypoint $p=p_{t}$ (or $p=\{p_{1},p_{2},...,p_{T}\}$), we aim to generate realistic images $I_{t}$ (or videos $V=\{I_{1},I_{2},...,I_{T}\}$) conditioned on $f,g,p$. The generated images (or videos) should be 1) faithful: the human attribute and background of the synthesis should be consistent with $f$ and $g$ from the reference image and the human subject should be aligned with the pose $p$; 2) generalizable: the model should be able to generalize to unseen humans, backgrounds and poses, without the need of human-specific fine-tuning; and 3) composable: the model should adapt to arbitrary composition of $f,g,p$ from different image/video sources to generate novel images/videos. In what follows, Section 3.1 briefly reviews the latent diffusion models and ControlNet, which are the basis of DisCo. Section 3.2 details the model architecture of DisCo with disentangled control of human foreground, background, and pose to enable faithful and fully composable human dance image/video synthesis. Section 3.3 presents how to further enhance the generalizability of DisCo, as well as the faithfulness in generated contents by pre-training human attributes from large-scale human images. The overview of DisCo can be found in Figure 1.
3.1 Preliminary
@@
Latent Diffusion Models (LDM) is a type of diffusion model that operates in the encoded latent space of an autoencoder $\mathcal{D}(\mathcal{E}(\cdot))$. An exemplary LDM is the popular Stable Diffusion (SD) [45] which consists an autoencoder VQ-VAE [62] and a time-conditioned U-Net [46] for noise estimation. A CLIP ViT-L/14 text encoder [43] is used to project the input text query into the text embedding condition $c_{\text{text}}$. During training, given an image $I$ and the text condition $c_{\text{text}}$, the image latent $z_{0}=\mathcal{E}(I)$ is diffused in $T$ time steps with a deterministic Gaussian process to produce the noisy latent $z_{T}\sim\mathcal{N}(0,1)$. SD is trained to learn the reverse denoising process with the following objective [45]:
@@
$L=\mathbb{E}_{\mathcal{E}(I),c_{\text{text}},\epsilon\sim\mathcal{N}(0,1),t}\left[\lVert\epsilon-\epsilon_{\theta}(z_{t},t,c_{\text{text}})\rVert_{2}^{2}\right],$ (1)
@@
where $t=\{1,...,T\}$, $\epsilon_{\theta}$ represents the trainable module. It contains a U-Net architecture composed of the convolution (ResBlock) and self-/cross-attention (TransBlock), which accepts the noisy latents $z_{t}$ and the text embedding condition $c_{\text{text}}$ as the input. After training, one can apply a deterministic sampling process (e.g., DDIM [56]) to generate $z_{0}$ and pass it to the decoder $\mathcal{D}$ towards the final image.
@@
ControlNet [71], built upon SD, manipulates the input to the intermediate layers of the U-Net in SD, for controllable image generation. Specifically, it creates a trainable copy of the U-Net down/middle blocks and adds an additional “zero convolution” layer. The outputs of each copy block is then added to the skip connections of the original U-Net. Apart from the text condition $c_{\text{text}}$, ControlNet is trained with an additional external condition vector $c$ which can be many types, such as edge map, depth map and segmentation.
@@ 3.2 Model Architecture with Disentangled Control
There are two typical control designs for the social media dance generation. The direct usage of ControlNet presents challenges due to the missing reference human image condition, which is critical for keeping the human identity and attributes consistent. Recent explorations in image variations [24] replace the CLIP text embedding with the image embedding as the condition, which can retain some high-level semantics from the reference image. Nevertheless, the geometric/structural control of the generated image is still missing. Meanwhile, simply combining these two designs does not work well in practice (Table 3 and 6).
@@
Figure 1: The proposed DisCo framework for social media human dance generation.
@@
To take the distinctive benefits of these two different control designs, we introduce a novel model architecture with disentangled control, to enable accurate alterations to the human pose, while simultaneously maintaining attribute and background stability. Meanwhile, it also facilitates full compositionality in the human dance synthesis (Figure LABEL:fig:intro_figb). Specifically, given a reference human image, we can first utilize an existing human matting method (e.g., SAM [28, 34]) to separate the human foreground from the background. Next, we explain how all three conditions, the human foreground $f$, the background $g$ and the desired pose $p$, are incorporated into DisCo.
@@
Referring Foreground via Cross Attention. To help model easily adapt to the CLIP image feature space, we first use the pre-trained image variation diffusion model [24] for the U-Net initialization. However, in contrast to using the global CLIP image embeddings employed by image variation methods, here we adopt the local CLIP image embeddings right before the global pooling layer, for more fine-grained human semantics encoding. Consequently, the original text embedding $c_{\text{text}}\in\mathbb{R}^{l\times d}$ is superseded by the local CLIP image embeddings of the human foreground $c_{f}\in\mathbb{R}^{hw\times d}$ to serve as the key and value feature in cross-attention layer, where $l,h,w,d$ represent the caption length, the height, width of the visual feature map and the feature dimension.
@@
Controlling Background and Pose via ControlNets. For pose $p$, we adopt the vanilla design of ControlNet. Specifically, we embed the pose image into the same latent space as the Unet input via four convolution layers, and dedicate a ControlNet branch $\tau_{\theta}$ to learn the pose control. For background $g$, we insert another ControlNet branch $\mu_{\theta}$ to the model. Notably, we propose to use the pre-trained VQ-VAE encoder $\mathcal{E}$ in SD, instead of four randomly initialized convolution layers, to convert the background image into dense feature maps to preserve intricate details. The remainder of the architecture for the background ControlNet branch follows the original ControlNet. As we replace the text condition with the referring foreground in the cross-attention modules, we also update the condition input to ControlNet as the local CLIP image feature of the referring foreground. As shown in Figure 0(a), the outputs of the two ControlNet branches are combined via addition and fed into the middle and up block of the U-Net.
@@
With the design of the disentangled controls above, we fine-tune DisCo with the same diffusion objective [45]:
@@
$\begin{split}L=&\mathbb{E}_{\mathcal{E}(I),c_{f},\tau_{\theta}(p),\mu_{\theta}(g),\epsilon\sim\mathcal{N}(0,1),t}[\lVert\epsilon\\ &\qquad\qquad-\epsilon_{\theta}(z_{t},t,c_{f},\tau_{\theta}(p),\mu_{\theta}(g))\rVert_{2}^{2}],\end{split}$ (2)
where $\epsilon_{\theta}$, $\tau_{\theta}$ and ${\mu_{\theta}}$ are the trainable network modules.
Figure 2: The detailed architecture of the ResBlock and TransBlock. The temporal module (dotted line) is optional.
@@
Specifically, $\epsilon_{\theta}$ represents the U-Net architecture composed of the convolution block (ResBlock) and self-/cross-attention block (TransBlock), which accepts the noisy latents $z_{t}$ and the referring foreground condition $c_{f}$ as the inputs. $\tau_{\theta}$ and ${\mu_{\theta}}$ represent the two ControlNet branches for pose condition $p$ and background condition $g$, respectively.
Temporal Modeling (TM). To achieve better temporal continuity for video output, we follow [54, 8] to introduce a 1D temporal convolution/attention layer after the existing 2D spatial ones in ResBlock and TransBlock (Figure 2). Notably, only the last yellow spatial attention module in TransBlock accept the CLIP image features for cross-attention operation. Besides, the temporal layers are zero-initialized, and connected via residual connections. Simultaneously, we adjust the model input from image to video by concatenating the pose $p$ from $n$ consecutive frames, duplicating the background $g$ and the initial noise for $n$ times.